Royal Company is considering the introduction of new equipment which cost $95,000. The equipment has useful life of four years and is in the three years property class for tax purposes. Shipping and installation charges are $5000. Machine has expected salvage value of $16500. No additional networking capital is needed. The new equipment will generate additional net operating cash flows before the depreciation and taxes as follows

Time Year 1 Year 2 Year 3 Year 4
Net Cash Flows 34000 37000 56000 35000

if the machine tax rate is 30%, calculate

1) The initial Cash flow
2) Interim incremental net cash flow
3) terminal year end incremental net cash flows (depreciation percentages 33.33%, 44.45%, 14.81% and 7.41%)
1
Expert's answer
2015-03-04T08:52:17-0500
Royal Company is considering the introduction of new equipment which cost $95,000. The equipment has useful life of 4 years and is in the three years property class for tax purposes. Shipping and installation charges are $5000, salvage value of $16500, tax rate is 30%.
Time Year 1 Year 2 Year 3 Year 4
NCF 34000 37000 56000 35000
1) The initial Cash flow is -$95,000.
2) Interim incremental net cash flow is -$5,000.
3) terminal year end incremental net cash flows (depreciation percentages 33.33%, 44.45%, 14.81% and 7.41%) are:
NCF1 = 34000*(1 - 0.3)*(1 - 0.3333) = 15867.46
NCF2 = 37000*0.7*(1 - 0.4445) = 14387.45
NCF3 = 56000*0.7*(1 - 0.1481) = 33394.48
NCF4 = 35000*(1 - 0.0741) = 32406.5